A boat moves with a constant velocity of 6m.s, starting from the position

shown. Find (a) the angle in order for the projectile to hit the boat 5sec

after starting under the condition given,(b) How high is the ill above the

water?


1
Expert's answer
2021-10-25T10:03:49-0400

The position of the boat and projectile must be the same in order for the projectile to hit the boat:


"6t+x_0=vt\cos\theta t,\\
\cos\theta= \frac{6\cdot5+x_0}{5v}."

The height:


"H=\frac{(v\sin\theta)^2}{2g}."